### v1

@@ Abstract
Markov state modeling has gained popularity in various scientific fields due to its ability to reduce complex time series data into transitions between a few states. Yet, current frameworks are limited by assuming a single Markov chain describes the data, and they suffer an inability to discern heterogeneities. As a solution, this paper proposes a variational expectation-maximization algorithm that identifies a mixture of Markov chains in a time-series data set. The method is agnostic to the definition of the Markov states, whether data-driven (e.g. by spectral clustering) or based on domain knowledge. Variational EM efficiently and organically identifies the number of Markov chains and dynamics of each chain without expensive model comparisons or posterior sampling. The approach is supported by a theoretical analysis and numerical experiments, including simulated and observational data sets based on Last.fm music listening, ultramarathon running, and gene expression. The results show the new algorithm is competitive with contemporary mixture modeling approaches and powerful in identifying meaningful heterogeneities in time series data.
@@ 1 Introduction
To address the complexity of high-dimensional data emerging from observations and experiments, a valuable strategy involves reducing data into a small collection of states. Often, the data comes in the form of dynamical trajectories which are modeled using a finite-state Markov chain that jumps between different regions (“states”) of phase space. This modeling approach is known as Markov state modeling [1], and it has become increasingly popular in chemistry [2, 3, 4], biology [5, 6, 7, 8], and climate science [9, 10, 11] over the past two decades.
In spite of the rapid uptake and development of Markov state modeling methods, there is a limiting deficit in strategies for handling heterogeneities that plague real data sets. This paper addresses the shortcoming in the Markov state modeling literature, as it makes the following argument:
Many time-series data sets are naturally modeled as mixtures of different Markov chains, rather than a single Markov chain.
The paper proposes that Markov state modeling, which is traditionally applied to one Markov chain at a time, can be applied to learn several Markov chains simultaneously. To make this approach computationally tractable, the paper presents a Markov chain mixture modeling algorithm that automatically detects the number of Markov chains and the dynamics of each chain. This method unlocks the ability to identify heterogeneities in data sets that existing approaches would miss. Indeed, in the numerical examples in Sec. 6, the differences between the Markov chains prove to be more enlightening than any single Markov chain on its own.
The proposed algorithm is based on variational expectation-maximization [12, Sec. 10], which is a well-established tool in the statistics literature. The goal of this paper is to communicate the usefulness of variational EM for Markov chain mixture analysis. The method can be applied to states that are defined through a data-driven approach (e.g., a clustering method) or by domain knowledge.
The adaptation of variational EM for the Markov state modeling context is novel. Variational EM was previously developed for Gaussian mixture models [13, 14, 12] and hidden Markov models [15, 16, 17], but apparently this is the first paper using the method for Markov chain mixtures.
As advantages, variational EM leads to a runtime that is linear in the number of trajectories $N$ and also linear in the number of model parameters. Additionally, variational EM enables automatic detection of the number of Markov chains [14, 18, 19]. Previously, researchers would fit several Markov chain mixture models with different numbers of components and then make time-consuming model comparisons using fit statistics such as AIC or BIC [20, 21, 22]. By eliminating the need for such model comparisons, variational EM can reduce the runtime to identify $k$ Markov chains by a factor of $k$ or more.
It is important to understand the practical and theoretical limitations of any new approach. As the first limitation, the Markov state mixture model is best suited for time-series data that can be plausibly modeled as (a) finite-state and (b) Markovian. This limitation is shared with the Markov state modeling more generally.
As the second limitation, variational EM does not always converge to the globally optimal parameters. Variational EM is fast enough that randomly initializing at $10^{2}$ different parameter values is a feasible strategy that leads to high-quality solutions. Yet, future research should investigate more principled initialization for variational EM.
As the last limitation, it is impossible for a Markov chain classification algorithm to achieve high accuracy when the chains are too short. This paper establishes a new result that the Kullback-Leibler divergence fundamentally limits the reliability of Markov chain mixture modeling (see Sec. 5). The Kullback-Leibler divergence increases linearly with the trajectory length $T$, so the optimal classification rate improves exponentially with $T$. Increasing the trajectory length is the single most important factor for boosting the classification accuracy of a Markov chain mixture model.
@@ 1.1 Plan for the paper
The rest of the paper is organized as follows. Section 2 introduces Markov state modeling and Markov state mixture modeling, Section 3 reviews previous related approaches, Section 4 introduces the variational EM algorithm, Section 5 provides theoretical analysis, Section 6 presents numerical experiments, and Section 7 concludes.
@@ 1.2 Notation
Scalars are in regular typeface: $k,s,t,T,N,D$. Lowercase letters $k,s,t$ indicate scalar quantities that can vary from line to line. Uppercase letters $T,N,D$ indicate scalar quantities that are fixed by the data set.
Vectors are typically written in bold lowercase typeface: $\bm{\mu},\bm{\nu},\bm{q}$. Matrices are written in bold uppercase typeface: $\bm{K},\bm{P}$. The entries of vectors and matrices are indicated with parentheses and regular typeface: $\mu(i)$, $D(i,i)$.
@@
Finite and infinite sets are in uppercase calligraphic letters, e.g., $\mathcal{X},\mathcal{Y}$. Elements of finite sets are often in lowercase Greek letters, e.g., $\alpha,\beta\in\mathcal{Y}$.
@@
2 Extending the Markov state model
A Markov state model is a classic approach for time series analysis with a history spanning centuries [1] and recent popularity in chemistry [2, 3, 4], biology [5, 6, 7, 8], and climate science [9, 10, 11]. A Markov state model approximates time series data as a finite-state Markov chain that jumps between suitably defined “states” of the system.
@@ 2.1 The finite-state assumption
There are various approaches for defining an appropriate set of finite states. As one approach, the states can be defined based on domain-area expertise. For example, user activity on MSNBC.com (“clickstreams”) can be reduced to a state space $\mathcal{Y}$ consisting of different types of webpages, such as the homepage (state 1), the news pages (state 2), and the sports pages (state 3) [22]. These state definitions are based on the MSNBC.com site map and the conventional division of articles into “sports” versus “news”.
As a more systematic alternative, the state space $\mathcal{X}$ can be decomposed using a distance function $d:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and a set of data centers $\{\bm{c}_{1},\ldots,\bm{c}_{s}\}$. This leads to the division into cells
@@
Any trajectory $\bm{X}$ on the state space $\mathcal{X}$ can be approximated by a trajectory $\bm{Y}$ on the reduced state space $\mathcal{Y}=\{1,\ldots,s\}$ by setting $Y_{t}=i$ if $X_{t}\in\mathcal{X}_{i}$ (with arbitrary tie-breaking at the boundaries).
@@
$k$-means clustering is common but not always effective at identifying finite states. Spectral clustering [24] is an alternative method that adapts to the nonlinear manifold structure of the input data. Spectral clustering can be applied to any state space (not just $\mathcal{X}=\mathbb{R}^{D}$), as long as there is a kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ which quantifies the similarity between data points. Fig.  1 illustrates spectral clustering on a spiral data set using the Gaussian kernel $k(\bm{x},\bm{y})=\exp(-\tfrac{1}{2\sigma^{2}}\lVert\bm{x}-\bm{y}\rVert^{2})$ with bandwidth $\sigma=1$. For a demonstration of spectral clustering on gene expression data, see Sec. 6.4.
@@
Algorithm 1 Spectral clustering [24]
1:Data points $\bm{x}_{1},\ldots,\bm{x}_{M}\in\mathcal{X}$, kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$, number of states $s$, embedding dimension $r$ (default $r=s$).
2:Clusters $\mathcal{X}=\mathcal{X}_{1}\cup\cdots\cup\mathcal{X}_{s}$.
3:Evaluate the kernel matrix $\bm{K}\in\mathbb{R}^{M\times M}$ with entries $K(i,j)=k(\bm{x}_{i},\bm{x}_{j})$.
4:Evaluate the diagonal matrix $\bm{D}\in\mathbb{R}^{M\times M}$ with entries $D(i,i)=\sum\nolimits_{j=1}^{M}K(i,j)$.
5:Calculate the dominant eigenvectors of $\bm{D}^{-1/2}\bm{K}\bm{D}^{-1/2}$, and set $\bm{V}=\begin{bmatrix}\bm{v}_{1}&\cdots&\bm{v}_{r}\end{bmatrix}$.
6:Update $\bm{V}\leftarrow\bm{D}^{-1/2}\bm{V}$.
7:Define the embedding $\bm{f}:\mathcal{X}\rightarrow\mathbb{R}^{r}$ via
$\bm{f}(\bm{x})=\sum\nolimits_{i=1}^{M}\bm{\alpha}(\cdot,i)k(\bm{x}_{i},\bm{x}),$
where $\bm{\alpha}\in\mathbb{R}^{r\times M}$ is chosen to solve $\bm{K}\bm{\alpha}^{T}=\bm{V}$. Note the embedded data points satisfy $\bm{f}(\bm{x}_{i})=\bm{V}(i,\cdot)^{T}$ for $1\leq i\leq M$.
8:Apply $k$-means to identify centers $\bm{c}_{1},\ldots,\bm{c}_{s}\in\mathbb{R}^{r}$ for the embedded data points.
9:Define clusters by
$\mathcal{X}_{i}=\bigl{\{}\bm{x}\colon\lVert\bm{f}(\bm{x})-\bm{c}_{i}\rVert\leq\lVert\bm{f}(\bm{x})-\bm{c}_{j}\rVert\text{ for }j\neq i\bigr{\}}.$
Alg. 1 provides simple, direct pseudocode for spectral clustering that emphasizes the cluster definitions. However, the pseudocode becomes prohibitively slow when the data size $M$ is large. For such cases, there is a faster spectral clustering implementation in [25] that reduces the cost from $\mathcal{O}(M^{2})$ operations to $\mathcal{O}(M)$ operations.
@@ 2.2 The Markovian assumption
$\mathbb{P}\{Y_{t}=\beta\,|\,Y_{0}=\alpha_{0},\ldots,Y_{t-1}=\alpha_{t-1}\}=P(\alpha,\beta),$
@@ 2.3 The one-chain assumption
The final assumption of a Markov state model is rarely stated yet fundamental. The standard presentations of Markov state modeling [28, 29, 1] all tacitly assume that a single Markov chain generates the trajectories. If the data set consists of multiple trajectories $\bm{Y}^{1},\bm{Y}^{2},\ldots,\bm{Y}^{n}$, the trajectories are all assumed to be generated from the same transition matrix $\bm{P}$.
The one-chain assumption is well-justified if the trajectories are computer simulations from a fixed model. Yet, the one-chain assumption becomes strained when faced with the heterogeneities of real-world observations. For example, is there really one type of web user on MSNBC.com? Is there really one type of commuter in Nanjing, China? Several papers in the statistics literature go beyond the one-chain assumption and instead they analyze MSNBC.com [22] users and Nanjing commuters [27] using a mixture of Markov chains with distinct transition matrices $\bm{P}_{1},\ldots,\bm{P}_{k}$. The resulting Markov chain mixture model will be described in the next section, and it will be the focus of the remainder of the paper.
@@ 2.4 The Markov chain mixture model
The Markov chain mixture model was presented in the papers [21, 30, 31], and it has been applied in areas including consumer behavior [22, 27], biology [32, 33], psychology [34], and economics [35]. However, the mixture model has apparently not been used in the Markov state modeling literature [28, 29, 1], which assumes a single Markov chain.
In the Markov chain mixture model, each trajectory $\bm{Y}$ is generated as follows:
@@
The variable $Z\in\{1,\ldots,k\}$ is a latent variable indicating the Markov chain label. $\mu(i)$ describes the probability of observing each label $Z=i$. The vectors $\bm{\nu}_{i}\in[0,1]^{s}$ and matrices $\bm{P}_{i}\in[0,1]^{s\times s}$ give the initialization probabilities and transition probabilities for the Markov chain with label $Z=i$. The total number of free parameters in the model is $ks^{2}-1$, and typical parameter values are $k=10^{0}$$10^{2}$ and $s=10^{0}$$10^{2}$.
3 Past algorithms for Markov chain mixtures
This section discusses past algorithms for fitting the parameters in a Markov chain mixture model, and it emphasizes two main themes. First, many researchers rely on the expectation-maximization (EM) algorithm for parameter inference (Sec. 3.1). Second, researchers have identified the key importance of the trajectory length $T$ for ensuring classification accuracy (Sec. 3.2).
@@ 3.1 The expectation-maximization algorithm
In several applications of Markov chain mixture modeling, the expectation-maximization (EM) algorithm gives fast and reliable parameter estimates [21, 30, 22, 27], especially when it is initialized carefully [36, 37, 38]. Alg. 2 provides a standard implementation of the EM algorithm for Markov chain mixtures. The algorithm iteratively refines the parameter estimates $\hat{\bm{\mu}}$, $\hat{\bm{\nu}}_{i}$, $\hat{\bm{P}}_{i}$, $\hat{Z}^{i}$, and outputs the converged values.
Algorithm 2 EM Algorithm [39]
1:Trajectories $(\bm{Y}^{n})_{1\leq n\leq N}$ and initial estimates ^$\hat{\mathbb{P}}\{Z^{n}=i\}$.
2:Refined parameter estimates $\hat{\bm{\mu}}$, $\hat{\bm{\nu}}_{i}$, $\hat{\bm{P}}_{i}$ and component labels $\hat{Z}^{n}\in\{1,\ldots,k\}$.
3:For each $n=1,\ldots,N$, form
$U^{n}(\alpha)$ $=\mathds{1}\{Y_{0}^{n}=\alpha\},$
$V^{n}(\alpha,\beta)$ $=\sum\nolimits_{t=0}^{T-1}\mathds{1}\{Y_{t}^{n}=\alpha,\,Y_{t+1}^{n}=\beta\},$
where $\mathds{1}\{\mathsf{S}\}$ is the function that takes the value 1 on set $\mathsf{S}$ and the value 0 otherwise.
4:for $t=1,2,\ldots$ do
5:     Update probabilities
$\hat{\mu}(i)$ $=\sum\nolimits_{n=1}^{N}\hat{\mathbb{P}}\{Z^{n}=i\},$
$\hat{\nu}_{i}(\alpha)$ $=\sum\nolimits_{n=1}^{N}\hat{\mathbb{P}}\{Z^{n}=i\}\,V^{n}(\alpha),$
$\hat{P}_{i}(\alpha,\beta)$ $=\sum\nolimits_{n=1}^{N}\,\hat{\mathbb{P}}\{Z^{n}=i\}V^{n}(\alpha,\beta).$
6:     Rescale probabilities so $\sum_{i=1}^{k}\hat{\mu}(i)=1$, $\sum_{\alpha=1}^{s}\hat{p}_{i}(\alpha)=1$, and $\sum_{\beta=1}^{s}\hat{p}_{i}(\alpha,\beta)=1$.
7:     Compute probabilities ^$\hat{\mathbb{P}}\{Z^{n}=i\}$ via
and choose $C_{n}$ so that $\sum_{i=1}^{k}\hat{\mathbb{P}}\{Z^{n}=i\}=1$.
8:     Calculate log likelihood $\mathcal{L}=\sum\nolimits_{n=1}^{N}\log C_{n}$.
9:     Break if $|\Delta\mathcal{L}|<10^{-12}\cdot NT$.
10:end for
11:Return the estimates $\{\hat{\bm{\mu}},\hat{\bm{\nu}}_{i},\hat{\bm{P}}_{i}\}$ and labels $\hat{Z}^{n}=\operatorname{argmax}_{1\leq i\leq k}\hat{\mathbb{P}}\{Z^{n}=i\}$.
As a limitation, EM is a nonconvex optimization algorithm that can converge to spurious local maxima in the likelihood landscape. Therefore, the accuracy of the EM parameter estimates depends on the initialization values, which can be chosen either randomly [22, 27] or by a more structured approach [21, 37, 38, 36]. In a random initialization, the user can generate probability estimates ^$\hat{\mathbb{P}}\{Z^{n}=i\}$ for $i=1,\ldots,k$ that are uniformly distributed on the probability simplex. A convenient procedure for generating uniformly distributed vectors on the simplex is to first produce $k$ independent exponential random variables and then normalize them to sum to one [40].
As another limitation, EM does not automatically determine the number of mixture components. Rather, the number of components must be specified as input, and determining the best number of components requires fitting various models and then performing a comparison using AIC or BIC [20, 21, 22]. This procedure can multiply the runtime by a factor of $k$ or more, assuming the user searches over $k$ different models. Sec. 4 will present an improvement of the EM algorithm called variational EM (VEM) that does automatically determine the number of components.

### v2

@@ Abstract
Markov state modeling has gained popularity in various scientific fields since it reduces complex time-series data sets into transitions between a few states. Yet common Markov state modeling frameworks assume a single Markov chain describes the data, so they suffer from an inability to discern heterogeneities. As an alternative, this paper models time-series data using a mixture of Markov chains, and it automatically determines the number of mixture components using the variational expectation-maximization algorithm. Variational EM simultaneously identifies the number of Markov chains and the dynamics of each chain without expensive model comparisons or posterior sampling. As a theoretical contribution, this paper identifies the natural limits of Markov state mixture modeling by proving a lower bound on the classification error. It then presents numerical experiments where variational EM achieves performance consistent with the theoretically optimal error scaling. The experiments are based on synthetic and observational data sets including Last.fm music listening, ultramarathon running, and gene expression. In each of the three data sets, variational EM leads to the identification of meaningful heterogeneities.
@@ 1 Introduction
This paper considers data emerging from observations, simulations, and experiments, which come in the form of dynamical trajectories. The trajectories can be modeled by a finite-state Markov chain that jumps between different regions (“states”) of phase space. This modeling approach is known as Markov state modeling [1], and it has become increasingly popular over the past two decades, including applications in chemistry [2, 3, 4], biology [5, 6, 7, 8], and climate science [9, 10, 11].
Markov state modeling has roots in classical statistical modeling, including clustering-based discretizations and finite-state Markov chains. Yet the data sets are changing. Observations, simulations, and experiments are now generating large collections of short, noisy, and heterogeneous trajectories that cannot be described by a single Markov chain. To address the contemporary modeling challenge, this work combines the discretization step in Markov state modeling with a Markov chain mixture approach. As specific contributions, the paper proposes an efficient modeling algorithm and it derives a non-asymptotic information-theoretic lower bound on the classification error achievable from finite trajectory lengths.
1.1 Goals and contributions
This paper argues that Markov state modeling, which is traditionally applied to one Markov chain at a time, can be applied to learn several Markov chains simultaneously. Markov state mixture modeling was previously proposed and developed in the works [12, 13, 14, 15, 16, 17, 18, 19, 20]. The paper builds on the existing research by asking and answering the following questions:
  • (a)
    What Markov chain mixture modeling approach can be easily adapted to high dimensions?
  • (b)
    What efficient algorithm can optimize the parameters in the model?
  • (c)
    What theoretical accuracy can the Markov mixture model attain?
For part (a), the paper applies a simple model [12] based on a mixture of Markov chains in a finite state space. Traditionally, the finite states were determined using expert knowledge of the predominant configurations of the system. However, the modern approach automatically identifies the states with a machine learning method such as spectral clustering.
For part (b), the paper adapts the variational expectation-maximization method [21, Sec. 10] to the setting of Markov chain mixture modeling. Variational EM has a major advantage over the classical EM algorithm, since classical EM requires fitting several models and then performing a post-hoc model comparison to determine the number of mixture components [22, 12, 15]. In contrast, variational EM determines the number of clusters automatically [23, 24, 25]. Here the variational EM approach is developed for finite-state Markov chains. Extending this framework to continuous-state or non-Markovian dynamics is a natural direction for future work.
For part (c), the paper quantifies the optimal theoretical accuracy of any Markov chain mixture model in terms of the Kullback-Leibler divergence between mixture components. The main theoretical contribution is the following theorem, with the proof appearing in Sec. 4.
Theorem 1 (Classification error bound).
Consider any statistical estimator $\hat{Z}$ for the label $Z$ in the Markov chain mixture model
$\mathbb{P}\{Z=i\}$ $=\mu(i),$
$\mathbb{P}\{Y_{0}=\alpha\,|\,Z=i\}$ $=\nu_{i}(\alpha),$
$\mathbb{P}\{Y_{t}=\beta\,|\,Z=i,\,Y_{t-1}=\alpha\}$ $=P_{i}(\alpha,\beta).$
The estimator can depend on the observed data $\bm{Y}$ and the model parameters but not on the unknown label $Z$. Then, the classification error is at least
In this expression, the Kullback-Leibler divergence is defined as
The summation runs over all possible trajectories of length $T$, and
$\mathbb{P}_{i}(\alpha_{0},\ldots,\alpha_{T})=\nu_{i}(\alpha_{0})\prod_{t=0}^{T-1}P_{i}(\alpha_{t},\alpha_{t+1}).$
The theorem lower bounds the classification error in terms of the Kullback–Leibler (KL) divergence between the trajectory distributions in the mixture model. Conceptually, it is a hypothesis-testing statement: if the mixture includes two similar trajectory distributions, no estimator can reliably assign trajectories to the correct distributions. The proof lower bounds the classification error by comparing pairs of components and quantifying how well they can be distinguished using $\operatorname{D}_{\rm KL}(\mathbb{P}_{i}\,\|\,\mathbb{P}_{j})$. The resulting bound exhibits the $e^{-\operatorname{D}_{\rm KL}}$ scaling familiar from the Bretagnolle–Huber inequality [26]; see Sec. 4 for further discussion.
In the context of Markov chain mixture modeling, the KL divergence in Theorem 1 is the divergence between length-$T$ trajectory distributions. Under mild regularity (e.g., ergodicity), this divergence grows linearly with $T$ at a rate determined by the stationary transition dynamics. Consequently, the lower bound in Theorem 1 decays exponentially in $T$, highlighting trajectory length as a dominant driver of identifiability.
@@ 1.2 Plan for the Paper
The rest of the paper is organized as follows. Section 2 introduces Markov state modeling and Markov state mixture modeling, Section 3 reviews previous related approaches, Section 3.3 introduces the variational EM algorithm, Section 4 provides theoretical analysis, Section 5 presents numerical experiments, and Section 6 concludes.
@@ 1.3 Notation
Scalars are in regular typeface: $k,s,t,T,N,D$. Lowercase letters $k,s,t$ indicate scalar quantities that can vary from line to line. Uppercase letters $T,N,D$ indicate scalar quantities that are fixed by the data set. Finite and infinite sets are in uppercase calligraphic letters, e.g., $\mathcal{X},\mathcal{Y}$. Elements of finite sets are often in lowercase Greek letters, e.g., $\alpha,\beta\in\mathcal{Y}$. Vectors are typically written in bold lowercase typeface: $\bm{\mu},\bm{\nu},\bm{q}$. Matrices are written in bold uppercase typeface: $\bm{K},\bm{P}$. The entries of vectors and matrices are indicated with parentheses and bold typeface: $\mu(i)$, $D(i,i)$.
@@
2 Modeling: Critiquing the Markov State Model
A Markov state model is a classic approach for time-series analysis with a long history [1]. It has recently become popular in chemistry [2, 3, 4], biology [5, 6, 7, 8], and climate science [9, 10, 11]. A Markov state model approximates time-series data as a finite-state Markov chain that jumps between suitably defined “states” of the system.
@@ 2.1 The Finite-state Assumption
There are various approaches for defining an appropriate set of finite states. As one approach, the states can be defined based on domain expertise. For example, user activity on MSNBC.com (“clickstreams”) can be reduced to a state space $\mathcal{Y}$ consisting of different types of webpages, such as the homepage (state 1), the news pages (state 2), and the sports pages (state 3) [15]. These state definitions are based on the MSNBC.com site map and the conventional division of articles into “sports” versus “news”.
As a more systematic alternative, the state space $\mathcal{X}$ can be decomposed using a distance function $d:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and a set of data centers $\{\bm{c}_{1},\ldots,\bm{c}_{s}\}$. This leads to the division into Voronoi cells
@@
Any trajectory $\bm{X}$ on the state space $\mathcal{X}$ can be approximated by a trajectory $\bm{Y}$ on the reduced state space $\mathcal{Y}=\{1,\ldots,s\}$ by setting $Y_{t}=i$ if $X_{t}\in\mathcal{X}_{i}$ with arbitrary tie-breaking at the boundaries. The rest of the paper will work with these discretized trajectories.
@@
$k$-means clustering is common but not always effective at identifying finite states. Spectral clustering [28] is an alternative method that adapts to the nonlinear manifold structure of the input data. Spectral clustering can be applied to any state space (not just $\mathcal{X}=\mathbb{R}^{D}$), as long as there is a kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ which quantifies the similarity between data points. Fig.  1 illustrates spectral clustering on a spiral data set using the Gaussian kernel $k(\bm{x},\bm{y})=\exp(-\tfrac{1}{2\sigma^{2}}\lVert\bm{x}-\bm{y}\rVert^{2})$ with bandwidth $\sigma=1$. Spectral clustering is now standard practice in Markov state modeling, so the algorithmic details are omitted; see [28] for the classical formulation and [29] for a scalable implementation based on randomized linear algebra. Later, Sec. 5.4 will apply spectral clustering to construct Markov states for gene expression data.
@@ 2.2 The Markovian Assumption
$\mathbb{P}\{Y_{t}=\beta\,|\,Y_{0}=\alpha_{0},\ldots,Y_{t-1}=\alpha_{t-1}\}$
$=P(\alpha_{t-1},\beta),$
@@
When strict Markovianity fails due to finite memory, a standard remedy is delay embedding: redefine the state at time $t$ as a tuple of lagged observations, e.g., $(Y_{t},Y_{t-1},\ldots,Y_{t-m+1})$. Delay embedding converts an $m$th-order Markov chain into a first-order Markov chain on an enlarged state space. Markov state modeling can be applied directly to the delay-embedded process, but the trade-off is an increased state-space size and hence more parameters to estimate.
@@ 2.3 The One-Chain Assumption
The final assumption of a Markov state model is rarely stated yet fundamental. The standard presentations of Markov state modeling [31, 32, 1] all assume that a single Markov chain generates the trajectories. If the data set consists of multiple trajectories $\bm{Y}^{1},\bm{Y}^{2},\ldots,\bm{Y}^{N}$, the trajectories are all assumed to be generated from the same transition matrix $\bm{P}$. The one-chain assumption is well-justified if the trajectories are computer simulations from a fixed model. Yet, the one-chain assumption is challenged by the heterogeneities of real-world observations and experiments. For example, is there really one type of web user on MSNBC.com? Is there really one type of commuter in Nanjing, China? Several papers in the statistics literature go beyond the one-chain assumption and instead they analyze MSNBC.com [15] users and Nanjing commuters [16] using a mixture of Markov chains with distinct transition matrices $\bm{P}_{1},\ldots,\bm{P}_{k}$.
The resulting Markov chain mixture model will be described in the next section, and it will be the focus of the remainder of the paper. Although joint inference of states and Markov chain mixture modeling parameters is possible [33], the paper assumes the state sequences have already been defined using domain knowledge or data-driven clustering. This design choice decouples state assignment from mixture inference to support scalable, domain-adaptable modeling.
@@ 2.4 The Markov chain mixture model
In a Markov chain mixture model, each trajectory $\bm{Y}$ is generated as follows:
@@
The variable $Z\in\{1,\ldots,k\}$ is a latent variable indicating the Markov chain label. $\mu(i)$ describes the probability of observing each label $Z=i$. The vectors $\bm{\nu}_{i}\in[0,1]^{s}$ and matrices $\bm{P}_{i}\in[0,1]^{s\times s}$ give the initialization probabilities and transition probabilities for the Markov chain with label $Z=i$. The total number of free parameters in the model is $ks^{2}-1$, and typical parameter values are $k=10^{0}$$10^{2}$ and $s=10^{0}$$10^{2}$. The Markov chain mixture model was presented in the papers [12, 13, 14], and it has been applied in areas including consumer behavior [15, 16], biology [17, 18], psychology [19], and economics [20].
3 Algorithms: Comparing Approaches for Markov Chain Mixtures
This section discusses past algorithms for fitting the parameters in a Markov chain mixture model, and it emphasizes three main themes. First, many researchers rely on the expectation-maximization (EM) algorithm for parameter inference (Sec. 3.1). Second, researchers have identified the key importance of the trajectory length $T$ for ensuring classification accuracy (Sec. 3.2). Third, this paper adapts the variational EM algorithm [21, Sec. 10] instead of traditional EM since it learns the number of mixture components automatically (Sec. 3.3).
@@ 3.1 The Expectation-Maximization Algorithm
In several applications of Markov chain mixture modeling, the expectation-maximization (EM) algorithm gives fast and reliable parameter estimates [12, 13, 15, 16], especially when it is initialized carefully [34, 35, 36]. In EM, each iteration alternates between (i) estimating conditional probabilities for the latent variables $\mathbb{P}\{Z^{n}=i\mid\bm{Y}^{n}\}$ and (ii) updating parameters $(\bm{\mu},\bm{\nu}_{i},\bm{P}_{i})$ using the expected counts of initial states and transitions under the latent variable distribution. Because EM for finite-state Markov-chain mixtures is standard (see, e.g., [37, 34]), this paper provides only a brief summary and focuses instead on the detailed exposition of variational EM in Sec. 3.3.
[…]
As another limitation, EM does not automatically determine the number of mixture components. Rather, the number of components must be specified as input. Many users fit several models with different numbers of components and then perform a comparison using AIC, BIC, or cross-validation [22, 12, 15]. When the user searches over a space of $k$ different models, this procedure can increase the runtime by a factor of $k$ or more. Sec. 3.3 will present an improvement of the EM algorithm called variational EM (VEM) that does automatically determine the number of components.